# Unified Domain Adaptive Semantic Segmentation

### v2

@@ III-B Video/Image Patch Templates Generation Across Domains
The domain gap becomes more evident when we observe diverse objects’ motion at the category-aware level. To mitigate the distinct point attributes across two domains existing in both video and image scenarios, we propose bridging and generalizing features by intra- and inter-domain mixing quad-directionally. This design necessitates the adaptive generation of reliable and diverse video/image patch sequences as templates, as denoted in Fig. 3. In this paper, we adaptively generate cross-domain patch templates online. Specifically, for training iteration , the source domain patch templates , and target domain patch templates , required for quad-directional mixing are derived from the previous iteration , where symbol “*” denotes templates. These patch templates contain pixel regions belonging to specific categories , which are randomly selected from segmentation label spaces and , respectively.
For UDA-VSS, video patch templates form a unified set of category-aware adjacent frames, labels/pseudo-labels, and optical flow. The target video patch template is defined as:
@@
where and corresponding to two selected categories, is the indicator function, refers to a pixel mapping operation by position-wise multiplication, and stands for a target filtering operation based on the confidence level of segmentation logits. Detailed depictions are shown in Fig. 4(a)(c).
@@
For UDA-ISS, the image patch template also forms a set but excludes temporal clues as in the video domain, which is denoted as for the target domain, and similarly for the source domain.
@@ III-C Intra-Mixed Video/Image Domain Adaptation
According to Fig. 1(a) and feature visualization in Fig. 7(a), despite the semantic similarity between source and target domains, a notable feature gap often persists in spatial structure and temporal coherence. Fig. 3(b) presents a depiction of four mixed paths of QuadMix. Rather than purely alternative training with cross-domain data [5, 37], or one way [19, 39, 23, 24]/bidirectional [21, 22] data pasting guided by offline-generated pseudo-labels via warm-up models stage-wisely, we propose a comprehensive QuadMix trained end-to-end. A more detailed process is shown in Fig. 4. In this section, we present the intra-mixed domain adaptation.
@@ III-C1 Target to Target Mixing
From the view of feature distribution, we consider source and target video/image domains as two distinct points featured with discontinuous and compact intra-domain feature distribution, and significant inter-domain feature gap. In this section, to address distinct point attributes within each domain for better gap bridging, we propose smoothing and generalizing intra-domains within an explicit feature space for the target domain.
For UDA-VSS, adaptive target video patch templates are iteratively integrated into target domain samples, guided by pseudo mask sequences. The process unfolds as follows:
for simplicity, Eq. (5) to Eq. (7) can be simplified as follows:
@@ III-C2 Source to Source Mixing
Discussion. Through intra-domain mixing, we adaptively generated intra-mixed source domain and intra-mixed target domain . As feature visualization in Fig. 7(b) for video examples, the feature distribution of the latter one is closer to the source domain, and both intra-mixed domains exhibit more generalized feature embeddings. Ablation studies in Table VI demonstrate that our intra-domain mixing outperforms purely alternating training, one-way or two-way mixing across domains. We attribute this to the enriched and continuous distribution of intra-mixed domains, which diminishes the distinct point attributes of the naive domains, thus effectively bridging the domain gap and facilitating knowledge transfer.
@@ III-D Inter-Mixed Video/Image Domain Adaptation
The feature gap bridging facilitated by intra-mixed domains prompts us to delve deeper into domain-wise interaction. In addition, existing methods [19, 21, 22, 23, 73, 24, 39] often focus purely on image-level mixing, neglecting the resulting semantic feature inconsistencies. Given our design of video/image patch templates at the pixel level, could we further enhance feature consistency across quad-mixed domains at the feature level? Driven by this, building upon our intra-mixed domains, we propose more enhanced inter-mixed domains and feature-level mixing, thereby bridging domain gaps comprehensively and alleviating feature inconsistencies.
@@ III-D1 Target to Intra-mixed Source Mixing
Following our design, inter-mixed domains should undergo feature-level interaction rather than being trained independently. To achieve this, we also utilize source and target patch templates to create more enhanced inter-mixed domains derived from our intra-domains, enabling the model to learn rich and invariant features across domains. The formula for the target to intra-mixed source domain mixing is as follows:
For UDA-ISS, the generated inter-mixed source domain sample is a set that integrates spatial elements. While for UDA-VSS, it further involves temporal elements, akin to Eq. (1), which encompasses video frames, labels, and inter-frame optical flow. Notably, the inter-mixed source domain integrates semantics from both source and target domains, featured with a more diverse and enhanced feature embedding space, which can facilitate the domain gap bridging.
@@ III-D2 Source to Intra-mixed Target Mixing
Compared to the large feature gap between naive source and target domains, the gap between the inter-mixed domains is effectively bridged, as depicted in Fig. 7(a)(e). This facilitates the transfer of rich knowledge learned from the source domain to the target domain, promoting effective domain adaptation. The partial labels from source domain patch templates within the inter-mixed target domain could aid in learning global features. Additionally, to enhance the diversity of patch templates from two domains, we ensure that categories of cross-domain templates are exclusive.
@@ III-D3 Feature-Level Template Mixing
Based on the design that the two inter-mixed domains encompass shared category-aware video/image templates from both the source and target domains, we encourage the model to alleviate feature inconsistencies arising from varying domain-specific semantic contexts and prioritize learning domain-invariant features of them.
For UDA-VSS, we adaptively perform feature-level template mixing across two inter-mixed domains to enhance the consistency learning of diverse spatio-temporal features, as depicted in Fig. 4(b). Firstly, the feature-level mask sequence is created to map spatio-temporal features from the inter-mixed source domain samples, and then the mapping results are fused with inter-mixed target domain features. The process is as follows:
@@
For UDA-ISS, corresponding to spatial image features, following a parallel manner.
Discussion. From a unified view, we propose a fairly simple, efficient, and practical solution to bridge video/image feature gaps and alleviate feature inconsistencies, facilitated by quad-mixing at pixel and feature levels. The diverse and rich domain-wise contexts aid in the domain-invariant feature learning, as illustrated in Fig. 1(d) and feature visualization in Fig. 7(e). Notably, we only utilize quad-mixed data and for training. Compared to incrementally using , our method yields superior results while reducing computational costs by a large margin. Detailed ablation analysis is provided in Table VI and Table XI.
@@ III-E Flow-Guided Feature Aggregation for Domain Alignment
To further mitigate domain gaps for UDA-VSS, we propose compressing video features into compact distribution for fine-grained alignment. Unlike only aligning spatial features of one single frame [39], we fully leverage temporal clues to generate cross-frame consistent target pseudo-labels, maintain temporal consistency of spatial features for each domain, and aggregate spatial and temporal features guided by optical flow for fine-grained domain alignment. As a result, aggregated video features from two domains exhibit an aligned and compact distribution, thus enhancing segmentation performance for the target domain. Note that it can also be extended to UDA-ISS.
@@ III-E1 Flow-Guided Spatio-Temporal Feature Extraction
In this paper, we employ a variant of ACCEL [57] as the basic architecture of video segmentation network, which involves two weight-shared CNN/Transformer-based segmentation branches and an optical flow branch. The segmentation branches capture the spatial feature, using two adjacent frames from the stacked as input. The optical flow is used to propagate temporal information across frames. We conduct feature-level template mixing proposed in Section III-D3 during forward training across inter-mixed domain samples.
For clarity, we illustrate with pure source and target domain in this part, showing that the inter-mixed source and target domain also adhere to the same paradigm. During source and target domain training, the features of adjacent frames are fused to generate holistic feature of frame , leveraging the video temporal continuity, where represents the deep feature of adjacent frames in this section. Specifically, we feed feature and optical flow-warped feature into a convolutional fusion layer to generate segmentation logits vector (abbreviated as ). The process is described as follows:
@@
For target pseudo-label generation, different from existing UDA-VSS works using a momentum network [21], [22], we employ a cross-frame propagation strategy to maintain the inherent temporal continuity. As shown in the left part of Fig. 2(a), consecutive input set from the target domain are fed into the segmentation network to generate the logits vector . Then, we yield pseudo-label by warping using optical flow , as described below:
@@ III-E2 Spatio-Temporal Feature Aggregation for Domain Alignment
Subsequently, we concatenate the aggregated video features along the category level for two domains to generate fine-grained video features, where the concatenation is denoted as . In this paper, we propose mitigating the fine-grained video feature gap using MMD (Maximum Mean Discrepancy) [29] minimization, benefiting from the optical flow-guided temporal information propagation and distribution alignment.
As for UDA-ISS, spatial aggregation can be used for fine-grained alignment without relying on temporal clues, i.e., optical flow guidance.
III-F Training Procedure and Objective Optimization
Our method follows a self-supervised learning paradigm trained end-to-end. The training algorithm is summarized in supplementary materials. The overall loss is defined as follows:
For UDA-VSS, specific losses are denoted as follows:
where denotes the video semantic segmentation network, means image enhancement, denotes cross-entropy loss, and and indicate two inter-mixed domains for simplicity.
For UDA-ISS, the model’s input in Eq. (22) and Eq. (24) should be replaced by a single image. Moreover, the sensitivity analysis for hyperparameters and will be specifically discussed in the experiments.
@@ III-F Training Procedure and Objective Optimization
| QuadMix (Trans.) | ✓ | 94.1 | 61.9 | 82.9 | 36.9 | 41.0 | 59.1 | 85.2 | 85.6 | 64.3 | 37.8 | 90.3 | 67.2 |
@@ IV-A Experimental Setups
IV-A1 UDA-VSS Benchmark
We conduct experiments on two challenging benchmarks: Synthia-Seq [8] Cityscapes-Seq [7] and VIPER [9] Cityscapes-Seq. Cityscapes-Seq (target domain) includes real city scene videos from 50 cities in Germany and neighboring countries. It comprises 2,975 training video clips and 500 validation video clips, and each clip contains continuous 30 frames. Synthia-Seq (source domain) contains 8,000 photo-realistic frames with dense labels. We use frames in the RGB folder and choose 11 categories for adaptation. VIPER (source domain) is another challenging synthetic video dataset, consisting of 133,670 frames of virtual videos rendered from urban landscape scenes in the virtual computer game “Grand Theft Auto V”. We use 13,367 frames with segmentation labels as another source domain, choosing 15 categories for adaptation. Our dataset setting is consistent with previous works [6, 37, 39, 38, 71, 36, 75], etc.
IV-A2 UDA-ISS Benchmark
@@
IV-A3 Implementation Details
Our model is implemented in PyTorch and trained on a single NVIDIA Tesla V100 GPU. The details are as follows: (i) As for the network, we employ a variant of ACCEL [57] for UDA-VSS, using DeepLabV2 [78] or MiT-B5 in DAformer [24] as the segmentation branch, and FlowNet [79] as the optical flow branch. For UDA-ISS, we only employ the segmentation branch. Notably, DeepLabV2 uses ResNet-101 [80] as its backbone, and both DeepLabV2 and MiT-B5 are pre-trained on ImageNet [81]; (ii) As for input resolution, for video tasks, we resize the resolution of Cityscapes-Seq, Synthia-Seq, and VIPER to , , and , respectively. While for image tasks, images undergo resizing to dimensions of , , and for Cityscapes, GTAV, and Synthia; (iii) We set batch size to 1 for video and 2 for image scenarios; (iv) For DeepLabV2, we use the SGD optimizer with a momentum of 0.9 and a weight decay of . The initial learning rate is for Synthia-Seq to Cityscapes-Seq and two image benchmarks, and for VIPER to Cityscapes-Seq. For transformer networks, we use an AdamW optimizer with a weight decay of 0.01 and betas set to (0.9, 0.999). The learning rate is initially set to for the encoder and for the decoder, with a linear warmup followed by linear decay; (v) Our models are trained for a maximum of 40k iterations for four benchmarks; (vi) Other issues: we apply Gaussian blur and color jitter to the target domain for image enhancement. Aligned with [6, 37, 39], we set video temporal deviation to 1. Additionally, the confidence threshold for target filtering operation is 0.9. During quad-directional mixing, source videos undergo random cropping to match the resolution of target videos. For image tasks, we set and to 1 for simplicity; (vii) In this paper, we use per-category IOU (Intersection over Union) and mIOU (mean IOU) as evaluation metrics. Experiment settings are consistent with prior works for fair comparison.
@@ IV-A3 Implementation Details
| QuadMix (Trans.) | ✓ | 87.3 | 43.8 | 87.3 | 25.2 | 40.0 | 36.9 | 86.7 | 20.8 | 90.3 | 65.8 | 86.8 | 48.6 | 65.6 | 37.6 | 49.7 | 58.2 |
IV-B Experimental Results
IV-B1 Comparison With State-of-the-Arts of UDA-VSS
In Table I and Table II, we compare our method with the state-of-the-art methods on Synthia-Seq Cityscapes-Seq and VIPER Cityscapes-Seq benchmark. To the best of our knowledge, recent methods [6, 71, 37, 36, 38, 75, 39] are the most superior works in UDA-VSS, all employing CNN architectures. In this paper, we also explore transformer-based networks for the first time. Moreover, we also compare several variants to UDA-VSS that are initially designed for UDA-ISS, including works [11, 12, 13, 14, 15, 16, 17, 18, 5, 19, 20]. Technically, following previous UDA-VSS works [6, 71, 37, 39], we replace their image segmentation backbones with a video segmentation network (denoted by †). Additionally, “Adv.” signifies methods that employ adversarial learning, and “SSL” denotes self-supervised learning. “Warm” denotes model parameter initialization using a warm-up model. “Trans.” denotes the transformer backbone, while others use ResNet-101 pre-trained on ImageNet.
Synthia-Seq Cityscapes-Seq. In Table I, we present experimental results comparing our method with state-of-the-art techniques on the Synthia-Seq Cityscapes-Seq. For clarity, the best per-category IOU and mIOU results are in bold, while suboptimal results are underlined. We observed that: (i) Our method is compatible with both CNN and transformer-based architectures. The proposed CNN-based Quadmix significantly outperforms the previous best CMOM [39] by 5.2 mIOU. When equipped with a transformer, we further boost mIOU by 5.7, achieving the highest result of 67.2; (ii) For CNN-based QuadMix, without a warm-up model, we achieved 59.6 mIOU, which is 4.3 higher than TPL-SFC [38] by a large margin. When using a warm-up model, we outperform the CMOM [39] by 5.2, which relies heavily on the warm-up model [6] for both model initialization and pseudo-label generation; (iii) Adversarial learning methods can improve the model’s adaptation but are less improving compared to self-supervised methods; (iv) Our method boosts the IOU for most categories, which can be attributed to our comprehensive quad-directional mixing, which effectively addresses distinct point attributes and feature inconsistencies, as well as flow-guided video feature aggregation for cross-domain alignment.
Viper Cityscapes-Seq. The domain gap between Viper and Cityscapes-Seq is much larger than that of Synthia-Seq and Cityscapes-Seq. As shown in Table II, our proposed QuadMix also achieves state-of-the-art results. We can conclude that: (i) For CNN-based QuadMix, we achieve 56.2 mIOU with a warm-up model and 54.9 mIoU without it; furthermore, we obtain 58.2 mIOU when we employ transformer-based network. All our results significantly outperform previous methods; (ii) Recent work I2VDA [71] focuses exclusively on transferring source spatial information to target domain videos, resulting in a target domain mIOU of 51.7. This is inferior to our model, which comprehensively investigates both spatial and temporal level domain shifts. (iii) Variants of image domain adaptation methods generally underperform on target videos compared to UDA-VSS methods, likely due to inadequate utilization of crucial temporal information.

### v3

@@ III-B Video/Image Patch Templates Generation Across Domains
The domain gap becomes more evident when we observe the motion of diverse objects at the category-aware level. To mitigate the intrinsic distributional gaps caused by intra-domain discontinuity and less generalizable feature distribution within both video and image scenarios, we propose generalizing features and bridging gaps by intra- and inter-domain mixing quad-directionally. This design necessitates the adaptive generation of diverse and reliable video/image patch sequences as templates, as denoted in Fig. 4. In this paper, we adaptively generate cross-domain patch templates online. Specifically, for training iteration , the source domain patch templates , and target domain patch templates , required for quad-directional mixing are derived from the previous iteration , where the symbol “*” denotes the templates. These patch templates contain pixel regions belonging to specific categories , which are randomly selected from segmentation label spaces and , respectively.
For video UDA-SS, patch templates form a unified set of category-aware adjacent frames, labels/pseudo-labels, and optical flow. The target video patch template is defined as:
@@
where and corresponding to two selected categories, is the indicator function, refers to the pixel mapping operation using Hadamard product, and stands for a target filtering operation based on the confidence level of segmentation logits. Detailed depictions are shown in Fig. 5(a)(c).
@@
For image UDA-SS, the image patch template also forms a set but excludes temporal cues, which is denoted as for the target domain, and similarly for the source domain. For convenience, we retain the subscript for images, considering them as a special case of videos.
@@ III-C Intra-Mixed Video/Image Domain Adaptation
According to Fig. 3(a) and feature visualization in Fig. 9(a), despite the semantic similarity between source and target domains, a notable feature gap often persists in spatial structure and temporal coherence. Fig. 4(b) presents a depiction of four mixed paths of QuadMix. Rather than purely alternative training with cross-domain data [10, 17], or one way [21, 2, 23, 1]/bidirectional [5, 22] data pasting guided by offline-generated pseudo-labels via warm-up models stage-wisely [2, 5], we propose a comprehensive QuadMix trained end-to-end. A more detailed process is shown in Fig. 5. In this section, we present the intra-mixed domain adaptation.
@@ III-C1 Target to Target Mixing
From the view of feature distribution, we consider source and target video/image domains as two distinct points featured with discontinuous intra-domain feature distribution, and significant inter-domain feature gap. To address the feature discontinuity within each domain for better gap bridging, we propose smoothing and generalizing intra-domains within an explicit feature space. In this part, we focus on the target domain.
For video UDA-SS, adaptive target video patch templates are iteratively integrated into target domain samples, guided by pseudo mask sequences. The process unfolds as follows:
for simplicity, Eq. (5) to Eq. (7) can be unified as follows:
@@ III-C2 Source to Source Mixing
Discussion. Through intra-domain mixing, we adaptively generate intra-mixed source and target domains. As feature visualization in Fig. 9(b) for video examples, the feature distribution of the latter one is closer to the source domain, and both intra-mixed domains exhibit more generalized feature embeddings. Ablation studies in Table VI demonstrate that our intra-domain mixing (Exp. ID 6) outperforms purely alternating training, one-way or two-way mixing across domains. We attribute this to the generalized and continuous distribution of intra-mixed domains, which diminishes the feature discontinuity within naive domains, thus effectively facilitating knowledge transfer.
@@ III-D Inter-Mixed Video/Image Domain Adaptation
The feature gap bridging facilitated by intra-mixed domains prompts us to delve deeper into domain-wise interaction. In addition, existing methods [21, 5, 22, 23, 3, 1, 2] often focus purely on image-level mixing, neglecting the resulting semantic feature inconsistencies, which will impact precise dense prediction on the target domain. Given our design of video or image patch templates at the pixel level, could we further enhance feature consistency across quad-mixed domains at the feature level? Driven by this, building upon our intra-mixed domains, we propose more enhanced inter-mixed domains and feature-level mixing, thereby bridging domain gaps comprehensively and alleviating feature inconsistencies.
@@ III-D1 Target to Intra-Mixed Source Mixing
Following our design, the mixed domains should undergo feature-level interaction rather than being trained in isolation. To achieve this, we also utilize source and target patch templates to create more enhanced inter-mixed domains derived from our intra-domains, enabling the model to learn rich and invariant features across domains. The formula for the target to intra-mixed source domain mixing is as follows:
For image UDA-SS, the generated inter-mixed source domain sample is a set that integrates spatial elements. While for video UDA-SS, it further involves temporal elements, akin to Eq. (1), which encompasses video frames, labels, and inter-frame optical flow. Notably, the inter-mixed source domain integrates semantics from both source and target domains, featured with a more diverse and enhanced feature embedding space, which can facilitate the domain gap bridging.
@@ III-D2 Source to Intra-Mixed Target Mixing
Compared to the large feature gap between naive source and target domains, the gap between the inter-mixed domains is narrowed effectively, as depicted in Fig. 9(a)(e). This facilitates the transfer of rich knowledge learned from the source domain to the target domain, promoting effective domain adaptation. The partial labels from source patch templates within the inter-mixed target domain could aid in the learning of global features. Notably, to enhance the diversity of patch templates, we ensure that categories of cross-domain templates are exclusive.
In this paper, we collectively refer to the two inter-mixed domains as quad-mixed domains because they are derived from intra-mixed domains and undergo quad-directional mixing.
@@ III-D3 Feature-Level Template Mixing
Based on the design that the generated quad-mixed domains encompass shared category-aware sample templates from both the source and target domains, we encourage the model to alleviate feature inconsistencies arising from domain-wise mixed-semantic contexts and prioritize learning domain-invariant features of them.
For video UDA-SS, to enhance the consistency learning of diverse spatio-temporal features, we perform feature-level template mixing across the generated quad-mixed domains adaptively, as depicted in Fig. 5(b). Specifically, feature-level mask sequence is generated to map spatio-temporal features from inter-mixed source domain samples, the mapping results are then fused with inter-mixed target domain features. The process is as follows:
@@
For image UDA-SS, corresponding to spatial image features, following a parallel manner.
Discussion. From a unified view, we propose a fairly simple, efficient, and practical solution to bridge video/image feature gaps and alleviate feature inconsistencies, facilitated by quad-directional mixing at pixel and feature levels. The diverse and rich mixed-domain contexts aid in the domain gap mitigation, as illustrated in Fig. 3(d) and feature visualization in Fig. 9(e).
@@ III-E Flow-Guided Feature Aggregation for Domain Alignment
To further mitigate domain gaps for video UDA-SS, we propose compressing video features into compact distribution for fine-grained alignment. Unlike only aligning spatial features of single frames [2], we fully leverage temporal cues to generate cross-frame consistent target pseudo-labels, maintain temporal consistency of spatial features for each domain, and aggregate spatio-temporal features guided by optical flow for precise domain alignment, thus enhancing segmentation for the target domain. Note that it is extendable to image UDA-SS.
@@ III-E1 Flow-Guided Spatio-Temporal Feature Extraction
In this paper, we employ a variant of ACCEL [37] as the video segmentation network, which involves two weight-shared CNN/transformer-based segmentation branches and an optical flow branch. The segmentation branches capture the spatial feature, taking stacked adjacent frames as the input. The optical flow is used to propagate temporal information across frames. We conduct feature-level template mixing proposed in Section III-D3 during forward training across quad-mixed domain samples.
For clarity, we illustrate with pure source and target domain, showing that the inter-mixed source and target domain also adhere to the same paradigm. During model training, the deep feature of stacked adjacent frames are fused to generate the holistic feature of frame , leveraging the video temporal continuity. Specifically, we feed feature and optical flow-warped feature into a convolutional fusion layer to generate segmentation logits vector (abbreviated as ). The process is described as follows:
@@
For target domain self-supervised training, we fully exploit inherent temporal continuity to generate cross-frame coherent pseudo-labels. As shown in the left part of Fig. 2(a), consecutive sets are fed into the segmentation network to generate logits vector . Then, we yield pseudo-label by warping using optical flow , as described below:
@@ III-E2 Spatio-Temporal Feature Aggregation for Domain Alignment
Subsequently, we concatenate the aggregated video features along the category level for two domains to generate fine-grained video features, where the concatenation is denoted as . We can mitigate the fine-grained video feature gap using MMD (Maximum Mean Discrepancy) [8] minimization, thanks to the optical flow-guided temporal information propagation and distribution alignment.
As for image UDA-SS, spatial feature aggregation can be used for fine-grained alignment without relying on temporal cues, i.e., optical flow guidance.
@@
| QuadMix (ViT) | ✓ | 94.1 | 61.9 | 82.9 | 36.9 | 41.0 | 59.1 | 85.2 | 85.6 | 64.3 | 37.8 | 90.3 | 67.2 |
III-F Model Training
The model is trained jointly using cross-domain samples, with the objective function defined as follows:
Each loss is discussed first for video UDA-SS as below.
where and indicate two quad-mixed domains for simplicity, denotes the semantic segmentation model, means image enhancement, and denotes cross-entropy loss.
where denotes the mathematical expectation, and represents the reproducing Hilbert space [61].
@@ IV-A Experimental Setups
IV-A1 Video UDA-SS Datasets
We conduct experiments on two challenging benchmarks: Synthia-Seq [62] Cityscapes-Seq [63] and VIPER [64] Cityscapes-Seq. Cityscapes-Seq (target domain) is a standard dataset for semantic urban scene understanding, featuring real-world videos from 50 cities in Germany and neighboring countries. It comprises 2,975 training video clips and 500 validation video clips, and each clip contains continuous 30 frames. Synthia-Seq (source domain) contains 8,000 photo-realistic frames with dense segmentation labels. We employ video frames in the RGB folder and choose 11 categories common with Cityscapes-Seq for adaptation. VIPER (source domain) is another standard synthetic video dataset, consisting of 133,670 frames of virtual videos rendered from urban landscape scenes in the virtual computer game “Grand Theft Auto V”. We use 13,367 frames with segmentation labels as another source domain, choosing 15 common categories with Cityscapes-Seq for adaptation. Notably, our dataset setting is consistent with previous works [16, 17, 2, 19, 44, 18, 20], etc.
IV-A2 Image UDA-SS Datasets
@@
| QuadMix (ViT) | ✓ | 87.3 | 43.8 | 87.3 | 25.2 | 40.0 | 36.9 | 86.7 | 20.8 | 90.3 | 65.8 | 86.8 | 48.6 | 65.6 | 37.6 | 49.7 | 58.2 |
IV-A3 Implementation Details
Our model is implemented in PyTorch and trained on a single NVIDIA Tesla V100 GPU. The details are as follows: (i) As for the network, we employ a variant of ACCEL [37] for video UDA-SS, using DeepLab-V2 [68] or MiT-B5 in DAFormer [1] as the segmentation branch, and FlowNet [69] as the optical flow branch. For image UDA-SS, we only employ the segmentation branch. Notably, DeepLab-V2 uses ResNet-101 [70] as its backbone, and both DeepLab-V2 and MiT-B5 are pre-trained on ImageNet [71]; (ii) As for input resolution, for video tasks, we resize the resolution of Cityscapes-Seq, Synthia-Seq, and VIPER to , , and , respectively. While for image tasks, images undergo resizing to dimensions of , , and for Cityscapes, GTAV, and SYNTHIA; (iii) We set batch size to 1 for video and 2 for image scenarios; (iv) For DeepLab-V2, we use the SGD optimizer with a momentum of 0.9 and a weight decay of . The initial learning rate is for Synthia-Seq to Cityscapes-Seq and two image benchmarks, and for VIPER to Cityscapes-Seq. For transformer networks, we use an AdamW optimizer with a weight decay of 0.01 and betas set to (0.9, 0.999). The learning rate is initially set to for the encoder and for the decoder, with a linear warmup followed by linear decay; (v) Our models are trained for a maximum of 40k iterations for four benchmarks; (vi) Other issues: we apply Gaussian blur and color jitter to the target domain for image enhancement. Aligned with [16, 17, 2], we set video temporal deviation to 1. Additionally, the confidence threshold for target filtering operation is 0.9. During quad-directional mixing, source videos undergo random cropping to match the resolution of target videos. For image tasks, we set and to 1 for simplicity; (vii) In this paper, we use per-category IOU (Intersection over Union) and mIOU (mean IOU) as evaluation metrics. Experiment settings are consistent with prior works for fair comparison.